OpenStudy (anonymous):

b+9/b-7 / 6b+54/b-1 divide and simplfy

OpenStudy (anonymous):

I couldn't understand exactly what you mean, i'm supposing it: \[\frac{\frac{b+9}{b+7}}{\frac{6b+54}{b-1}}\] \[\frac{b+9}{6.(b+9)}.\frac{b-1}{b+7}=\frac{b-1}{6b+42}\]

OpenStudy (anonymous):

the answer is (b-1)/6(b-7)

jhonyy9 (jhonyy9):

(b+9)/(b-7)/(6b+54)/(b-1) =(b+9)/(b-7)x(b-1)/(6(b+9)) = (b-1)/(6(b-7))